OpenStudy (anonymous):

Marie sells magazine subscriptions. She earns $15 a week plus $3 for each subscription she sells. How many subscriptions must she sell to earn $90 in one week?

Parth (parthkohli):

=> 15 + 3x = 90 => 3x = 75 => x = 75/3 = 25 So, she has to sell 25 subscriptions to earn 90 dollars in one week.